if XYZ invests 3,100 today and 3,100 in 1 year in an account that has an expected annual return of 16.8 percent, compounded quarterly, then how much money will she have in her account in 3 years?
Answer #1

Future value = Present value *(1+rate)period

rate =16.8*3/12 i.e.4.2% qrtly

Period =3*4 i.e 12 for 1st contribution and 2*4 i.e.8 for 2nd

Future value =3100(1+4.2%)12+3100(1+4.2%)8 i.e. $9387.23 will have in year 3